What is the mass of one molecule of acetamiprid
1
Expert's answer
2019-10-14T14:29:48-0400

Acetamiprid is an organic compound with the chemical formula C10H11ClN4.

Mr(C10H11ClN4) = 222.5.

m(C10H11ClN4) = Mr(C10H11ClN4) × 1u = 222.5 × 1.66 × 10-27 kg = 369.35×10-27 kg ≈ 3.69×10-25kg.